# Reflected Brownian Motion in a wedge: sum-of-exponential absorption probability at the vertex and differential properties

### v1

@@ Abstract
We study a Brownian motion with drift in a wedge of angle $\beta$ which is obliquely reflected on each edge along angles $\varepsilon$ and $\delta$. We assume that the classical parameter $\alpha=\frac{\delta+\varepsilon-\pi}{\beta}$ is greater than $1$ and we focus on transient cases where the process can either be absorbed in the vertex or escape to infinity. We show that $\alpha\in\mathbb{N^{*}}$ is a necessary and sufficient condition for the absorption probability, seen as a function of the starting point, to be written as a finite sum of terms of exponential product form. In such cases, we give expressions for the absorption probability and its Laplace transform. When $\alpha\in\mathbb{Z}+\frac{\pi}{\beta}\mathbb{Z}$ we find an explicit D-algebraic expression for the Laplace transform. Our results rely on Tutte’s invariant method and a recursive compensation approach.
@@ 1 Introduction
In dimension one, it is known that a standard Brownian motion with positive drift $\mu>0$ started at $u>0$ has probability $e^{-2\mu u}$ to reach 00. A simple way of achieving this result is to use Girsanov’s theorem and the reflection principle. In dimension 2, we consider an obliquely reflected Brownian motion in a cone with drift belonging to the interior of the cone and directions of reflection strongly oriented towards the apex of the cone. A phenomenon of competition between the reflections and the drift appears and the process is either absorbed at the vertex or escapes to infinity. Lakner, Liu, and Reed [16] studied this absorption phenomenon and showed the existence and uniqueness of a solution to the absorbed process. Ernst et al. [7] were able to obtain a general formula for the probability of absorption at the vertex using Carleman’s boundary value problems theory. In particular, they characterised the cases where this probability has an exponential product form, i.e. when the reflection vectors are opposite. Franceschi and Raschel [10] then generalised this result to higher dimensions by showing that the coplanarity of the reflection vectors was a necessary and sufficient condition for the absorption probability to have an exponential product form. In a sense, this condition can be seen as dual to the classical skew symmetry condition discovered by Harrison and Williams [12, 14, 22] which characterises cases where the stationary distribution is exponential. In dimension 2, when the process is recurrent, Dieker and Moriarty [6], preceded by Foshini [11] in the symmetric case, determined a necessary and sufficient condition for the stationary distribution to be a sum of exponentials terms of product form. It is therefore very natural to look for an analogous result to the one of Dieker and Moriarty. This article aims to find, when the process is transient, a necessary and sufficient condition for the absorption probability to be a sum-of-exponentials function of the starting point and to compute this probability. We also identify other remarkable cases where the Laplace transform of the absorption probability is D-algebraic.
@@
which is a famous key parameter in the SRBM literature. As a general rule, such a process is most of the time studied in the literature in the case where $\alpha<1$, i.e. in the case where the process is a semimartingale markov process, see the seminal work of Varadhan and Williams [19, 21]. We will not give here a precise mathematical definition of the process, which can be found in many articles, see the survey of Williams [23]. We will simply point out that it behaves like a standard Brownian motion with drift inside the cone, it is reflected in a given direction when it touches an edge (being pushed by the local time on the boundary) and it spends zero time at the vertex of the cone. The famous skew symmetric condition, where the stationary distribution has an exponential product form, corresponds to $\alpha=0$, and Dieker and Moriarty’s condition for a sum-of-exponential stationary density corresponds to $\alpha\in-\mathbb{N}$. The dual skew symmetric case [7, 10], where the escape probability has an exponential product form, correspond to $\alpha=1$. For our purposes, in this article, we will assume that
@@
$\alpha\in\mathbb{N}^{*}$ (3)
plus the small technical condition
@@
In Theorem 11 we state another more general and stronger result which explicitly determines the Laplace transform of the absorption probability in terms of a conformal gluing function when
@@
In this case, we also find the differential nature of the Laplace transform. In other words we find sufficient conditions on $\alpha$ for the Laplace transform to be rational, algebraic (i.e. satisfying a polynomial equation with coefficients in the field of rational functions), D-finite (i.e. satisfying a linear differential equation with coefficients in the field of rational functions) or D-algebraic (i.e. satisfying a polynomial differential equation with coefficients in the field of rational functions). The differential nature of the Laplace transform reflects in various ways on the absorption probability itself. For example, if it is rational it implies that the absorption probability is a linear combination of exponentials multiplied by polynomials. If it is D-algebraic it will give a recurrence relation for the moments. We refer to the introduction of [3] which explains in more detail the interest of such a classification in this hierarchy of functions:
@@
$\alpha\in\mathbb{N}^{*}$ $\pi/\beta\in\mathbb{Q}$ and $\alpha\in\mathbb{Z}+\frac{\pi}{\beta}\mathbb{Z}$ $\alpha\in\mathbb{N}^{*}+\frac{\pi}{\beta}\mathbb{Z}$ $\alpha\in\mathbb{Z}+\frac{\pi}{\beta}\mathbb{Z}$
@@
Section 2 presents the preliminaries needed to prove our results. For technical reasons, we first transfer the problem initially defined in a wedge into a quadrant thanks to a simple linear transform. The starting point of the proof is a kernel functional equation satisfied by the Laplace transform of the absorption probability as a function of the starting point. This equation is derived from a partial differential equation solved by the probability of absorption. This functional equation leads to a boundary value problem (BVP) already studied in [7]. In Section 3, we apply successfully Tutte’s invariant method [20] to this BVP finding some decoupling functions, in a similar way to what was done in the recurrent case for the stationary distribution [3, 9]. We then compute explicitly the Laplace transform, see Theorems 9 and 11. Inverting the bivariate Laplace Transform is no easy task because of a complicated factorization of a two variable polynomial by the kernel. In Section 4, we then offer a geometrical way to construct the solutions inspired by the compensation approach developed with success in the discrete case for some queueing problems and random walks by Adan, Wessels, and Zijm [1].
@@
It is also possible to interpret our problem as the study of the probability of triple collisions for transient competing particle systems with asymmetric collisions. Indeed, a Brownian motion reflected in a quadrant is nothing more than the gap process of such a system made of three particles, and reaching the vertex of the quadrant is equivalent to a triple collision. A very interesting literature is devoted to the study of the absence or presence of such collisions, and as we cannot claim exhaustiveness in these few lines we will limit ourselves to mentioning the articles by Ichiba, Karatzas, and Shkolnikov [15] and Bruggeman and Sarantsev [4].
@@ 2 Preliminaries
The results stated in the introduction for the standard Brownian motion $\widetilde{Z}_{t}$ reflected in a cone $C$ of angle $\beta\in(0,\pi)$, drift angle $\theta$ and reflection angles $\delta$ and $\varepsilon$ will be proved by considering a Brownian motion $Z_{t}$ reflected in the quadrant $\mathbb{R}_{+}^{2}$ with a covariance matrix, a drift and a reflection matrix noted respectively
@@
which maps the cone $C$ onto the first quadrant $\mathbb{R}_{+}^{2}=\phi(C)$, and $\widetilde{Z}_{t}$ onto $Z_{t}=\phi(\widetilde{Z}_{t})$.
We have of course $\phi(0)=0$ and $T=\inf\{t>0:\widetilde{Z}_{t}=0\}=\inf\{t>0:Z_{t}=0\}$. It is then equivalent to compute the absorption probability for the process $\widetilde{Z}_{t}$ starting from $(\widetilde{u},\widetilde{v})\in C$ and for the process $Z_{t}$ starting from $(u,v)=\phi(\widetilde{u},\widetilde{v})\in\mathbb{R}_{+}^{2}$. We denote the escape probability
@@
The set $\mathcal{S}$ is an elliptic curve that passes through the origin. For $x\in\mathbb{C}$, the equation $K(x,y)=0$ in $y$ is quadratic, and has therefore two solutions $Y^{+}(x)$ and $Y^{-}(x)$ in $\mathbb{C}$:
@@
$x^{\pm}:=\frac{(\mu_{2}\sigma_{12}-\mu_{1}\sigma_{22})\pm\sqrt{(\mu_{2}\sigma_{12}-\mu_{1}\sigma_{22})^{2}+\det(\Sigma)\mu_{2}^{2}}}{\det(\Sigma)}.$
The branches $X^{+}$ and $X^{-}$ are analytic on $\mathbb{C}\setminus((-\infty,y^{-}]\cup[y^{+},+\infty))$, and $Y^{+}$ and $Y^{-}$ are analytic on $\mathbb{C}\setminus((-\infty,x^{-}]\cup[x^{+},+\infty))$. It will be handy to work with the following rational uniformization of $\mathcal{S}$, first stated in [8, Proposition 5], defined by
@@
Figure 3: On the left, the points introduced in Lemmas 20 and 13 on the unit complex circle $\mathbb{U}$; on the right, their images by $(\bm{\mathrm{x}},\bm{\mathrm{y}})$ on the ellipse $\mathcal{S}\cap\mathbb{R}^{2}$. For the sake of readability, points on the ellipse are labelled with their preimages. Invariances in (19) are represented by the dotted arrows.
@@
    $G(y)=\frac{k_{1}}{k_{2}}(X^{+}(y),y)\frac{k_{2}}{k_{1}}(X^{+}(y),\overline{y}).$ (26)
@@ 3.1 Decoupling and Tutte’s invariant
The same lemma gives, for some constant $\hat{C}$, the asymptotics
@@ 3.2 Explicit expression for the Laplace transforms
We now state our first main result when $\alpha\in\mathbb{N}^{*}$.
Theorem 9 (Laplace transforms, $\alpha\in\mathbb{N}^{*}$).
If $\alpha\in\mathbb{N}^{*}$ then the rational function defined by
$L(x,y):=\frac{k_{1}(x,y)P(x)+k_{2}(x,y)Q(y)}{K(x,y)}$
@@
Assuming that $\alpha=d-1\in\mathbb{N}^{*}$, the polynomials $P$ and $Q$ given in (31) form a decoupling pair by Proposition 7. The boundary value problem of Proposition 3 thus implies that $Q\varphi_{1}$ is an invariant, see (28). According to Lemma 6, we only need to prove that $Q\varphi_{1}$ doesn’t have any pole on $\overline{\mathcal{G}}_{\mathcal{H}}$, and has a finite limit as $y$ goes to +$\infty$+ ∞. By (24) and (31), and since $\alpha=d-1$ we have
@@
Comparing with Definition 4 we can see that the constant $\lambda$ given in (37) is equal to $-1$. The polynomial $k_{2}Q+k_{1}P$ vanishes on $\mathcal{S}$ the set of the zeros of $K$. By Hilbert’s Nullstellensatz,
@@
where $(K):=\{LK,L\in\mathbb{C}[X,Y]\}\text{ and }\sqrt{(K)}:=\{H\in\mathbb{C}[X,Y]:\exists m\in\mathbb{N}\text{ s.t. }H^{m}\in(K)\}$ are the ideal generated by $K$ and its radical. Here, $K$ is irreducible which implies that $\sqrt{(K)}=(K)$. This shows that there exists $L\in\mathbb{C}[X,Y]$ such that $L(x,y)K(x,y)=k_{1}(x,y)P(x)+k_{2}(x,y)Q(y).$ We see that $L(x,y)\in\mathbb{R}$ for all $(x,y)\in\mathbb{R}^{2}$, so the coefficients of $L$ must be real. Substituting the values for $\varphi_{1}$ and $\varphi_{2}$ into the functional equation (13) yields $\varphi(x,y)=L(x,y)/(P(x)Q(y))$. ∎
@@
[…]
We deduce sufficient conditions for $\varphi_{1}$, $\varphi_{2}$ and $\varphi$ to belongs to the hierarchy (6). If $\alpha\in\mathbb{N}^{*}$ these Laplace transforms are rational, if $\pi/\beta\in\mathbb{Q}$ and $\alpha\in\mathbb{Z}+\frac{\pi}{\beta}\mathbb{Z}$ they are algebraic, if $\alpha\in\mathbb{N}^{*}+\frac{\pi}{\beta}\mathbb{Z}$ they are D-finite and if $\alpha\in\mathbb{Z}+\frac{\pi}{\beta}\mathbb{Z}$ they are D-algebraic.

### v2

@@ Abstract
We study a Brownian motion with drift in a wedge of angle $\beta$ which is obliquely reflected on each edge along angles $\varepsilon$ and $\delta$. We assume that the classical parameter $\alpha=\frac{\delta+\varepsilon-\pi}{\beta}$ is greater than $1$ and we focus on transient cases where the process can either be absorbed at the vertex or escape to infinity. We show that $\alpha\in\mathbb{N}$ is a necessary and sufficient condition for the absorption probability, seen as a function of the starting point, to be written as a finite sum of terms of exponential product form. In such cases, we give expressions for the absorption probability and its Laplace transform. When $\alpha\in\mathbb{Z}+\frac{\pi}{\beta}\mathbb{Z}$ we find an explicit differentially-algebraic expression for the Laplace transform. Our results rely on Tutte’s invariant method and a recursive compensation approach.
@@ 1 Introduction
In dimension one, it is known that a standard Brownian motion with positive drift $\mu>0$ started at $u>0$ has probability $e^{-2\mu u}$ to reach 00. A simple way of achieving this result is to use Girsanov’s theorem and the reflection principle. In dimension 2, we consider an obliquely reflected Brownian motion in a cone with drift belonging to the interior of the cone and directions of reflection strongly oriented towards the apex of the cone. A phenomenon of competition between the reflections and the drift appears and the process is either absorbed at the vertex or escapes to infinity. Lakner, Liu, and Reed [20] studied this absorption phenomenon and showed the existence and uniqueness of a solution to the absorbed process. Ernst et al. [10] were able to obtain a general formula for the probability of absorption at the vertex using Carleman’s boundary value problems theory. In particular, they characterised the cases where this probability has an exponential product form, i.e. when the reflection vectors are opposite. Franceschi and Raschel [13] then generalised this result to higher dimensions by showing that the coplanarity of the reflection vectors was a necessary and sufficient condition for the absorption probability to have an exponential product form. In a sense, this condition can be seen as dual to the classical skew symmetry condition discovered by Harrison and Williams [15, 17, 29] which characterises cases where the stationary distribution is exponential. In dimension 2, when the process is recurrent, Dieker and Moriarty [9], preceded by Foshini [14] in the symmetric case, determined a necessary and sufficient condition for the stationary distribution to be a sum of exponentials terms of product form. It is therefore very natural to look for an analogous result to the one of Dieker and Moriarty. This article aims to find, when the process is transient, a necessary and sufficient condition for the absorption probability to be a sum-of-exponentials function of the starting point and to compute this probability. We also identify other remarkable cases where the Laplace transform of the absorption probability is differentially-algebraic (D-algebraic), i.e. solution of a polynomial equation in the function, its derivatives, and the independent variables, with coefficients in $\mathbb{R}$.
@@
which is a famous key parameter in the SRBM literature. As a general rule, such a process is most of the time studied in the literature in the case where $\alpha<1$, i.e. in the case where the process is a semimartingale markov process, see the seminal work of Varadhan and Williams [26, 28]. We will not give here a precise mathematical definition of the process, which can be found in many articles, see the survey of Williams [30]. We will simply point out that it behaves like a standard Brownian motion with drift inside the cone, it is reflected in a given direction when it touches an edge (being pushed by the local time on the boundary) and it spends zero time at the vertex of the cone. The famous skew symmetric condition, where the stationary distribution has an exponential product form, corresponds to $\alpha=0$, and Dieker and Moriarty’s condition for a sum-of-exponential stationary density corresponds to $\alpha\in-\mathbb{N}\cup\{0\}$. The dual skew symmetric case [10, 13], where the escape probability has an exponential product form, correspond to $\alpha=1$. For our purposes, in this article, we will assume that
@@
$\alpha\in\mathbb{N}:=\{1,2,3,\dots\}$ (3)
plus the condition
@@
In Theorem 12 we state another more general and stronger result which explicitly determines the Laplace transform of the absorption probability in terms of a Gauss hypergeometric function when
@@
In this case, we also find the differential nature of the Laplace transform. In other words we find sufficient conditions on $\alpha$ for the Laplace transform to be rational, algebraic (i.e. satisfying a polynomial equation with coefficients in the field of rational functions  over $\mathbb{R}$), D-finite (i.e. satisfying a linear differential equation with coefficients in the field of rational functions  over $\mathbb{R}$) or D-algebraic (i.e. satisfying a polynomial differential equation with coefficients in the field of rational functions  over $\mathbb{R}$). The differential nature of the Laplace transform reflects in various ways on the absorption probability itself. For example, if it is rational it implies that the absorption probability is a linear combination of exponentials multiplied by polynomials. If it is D-algebraic it will give a recurrence relation for the moments. We refer to the introduction of [5] which explains in more detail the interest of such a classification in this hierarchy of functions:
@@
$\alpha\in\mathbb{N}$ $\pi/\beta\in\mathbb{Q}$ and $\alpha\in\mathbb{Z}+\frac{\pi}{\beta}\mathbb{Z}$ $\alpha\in\mathbb{N}+\frac{\pi}{\beta}\mathbb{Z}$ $\alpha\in\mathbb{Z}+\frac{\pi}{\beta}\mathbb{Z}$
@@
Section 2 presents the preliminaries needed to prove our results. For technical reasons, we first transfer the problem initially defined in a wedge into a quadrant thanks to a simple linear transform. The starting point of the proof is a kernel functional equation satisfied by the Laplace transform of the absorption probability as a function of the starting point. This equation is derived from a partial differential equation solved by the probability of absorption. This functional equation leads to a boundary value problem (BVP) already studied in [10]. In Section 3, we apply successfully Tutte’s invariant method [25] to this BVP finding some decoupling functions, in a similar way to what was done in the recurrent case for the stationary distribution [5, 12]. We then compute explicitly the Laplace transform, see Theorems 10 and 12. Inverting the bivariate Laplace Transform is no easy task because of a complicated factorization of a two variable polynomial by the kernel. In Section 4, we then offer a geometrical way to construct the solutions inspired by the compensation approach developed with success in the discrete case for some queueing problems and random walks by Adan, Wessels, and Zijm [1].
@@
It is also possible to interpret our problem as the study of the probability of triple collisions for transient competing particle systems with asymmetric collisions. Indeed, a Brownian motion reflected in a quadrant is nothing more than the gap process of such a system made of three particles, and reaching the vertex of the quadrant is equivalent to a triple collision. A very interesting literature is devoted to the study of the absence or presence of such collisions, and as we cannot claim exhaustiveness in these few lines we will limit ourselves to mentioning the articles by Ichiba, Karatzas, and Shkolnikov [19], Bruggeman and Sarantsev [7], and Sarantsev [23]. In stochastic finance, Banner, Fernholz and Karatzas [3] shown strong connections between rank-based models (such as Atlas models) and the reflected Brownian motion.
In queueing theory, the reflected Brownian motion can be thought as the (scaled) limit of the queueing length process. Such convergence results are refered to as heavy traffic limit theorems, see the founding article by Harrison [15] and the classic book by Whitt [27]. This is still a very active research fields with many applications, see for example [2, 6].
@@ 2 Preliminaries
The results presented in this paper are particularly neat when expressed in terms of the angles that define the process in a cone, whereas the proofs are simpler for a process in the quadrant. This is why the results stated in the introduction for the standard Brownian motion $\widetilde{Z}_{t}$ reflected in a cone $C$ of angle $\beta\in(0,\pi)$, drift angle $\theta$ and reflection angles $\delta$ and $\varepsilon$ will be proved by considering a Brownian motion $Z_{t}$ reflected in the quadrant $\mathbb{R}_{+}^{2}$ with a positive-definite covariance matrix, a drift and a reflection matrix noted respectively
@@
which maps the cone $C$ onto the first quadrant $\mathbb{R}_{+}^{2}=\phi(C)$, and $\widetilde{Z}_{t}$ onto $Z_{t}=\phi(\widetilde{Z}_{t})$. We have of course $\phi(0)=0$ and $T=\inf\{t>0:\widetilde{Z}_{t}=0\}=\inf\{t>0:Z_{t}=0\}$. It is then equivalent to compute the absorption probability for the process $\widetilde{Z}_{t}$ starting from $(\widetilde{u},\widetilde{v})\in C$ and for the process $Z_{t}$ starting from $(u,v)=\phi(\widetilde{u},\widetilde{v})\in\mathbb{R}_{+}^{2}$. We denote the escape probability
@@
Note that thanks to this linear transformation, our results generalise immediately to all Brownian motions with any covariance matrix (possibly different from the identity) in any convex cone.
@@
For $x\in\mathbb{C}$, the equation $K(x,y)=0$ in $y$ is quadratic, and has therefore two solutions $Y^{+}(x)$ and $Y^{-}(x)$ in $\mathbb{C}$:
@@
$x^{\pm}:=\frac{(\mu_{2}\sigma_{12}-\mu_{1}\sigma_{22})\pm\sqrt{(\mu_{2}\sigma_{12}-\mu_{1}\sigma_{22})^{2}+\det(\Sigma)\mu_{2}^{2}}}{\det(\Sigma)}.$ (17)
The branches $X^{+}$ and $X^{-}$ are analytic on $\mathbb{C}\setminus((-\infty,y^{-}]\cup[y^{+},+\infty))$, and $Y^{+}$ and $Y^{-}$ are analytic on $\mathbb{C}\setminus((-\infty,x^{-}]\cup[x^{+},+\infty))$.
Lemma 2.
The kernel $K$ is irreducible over $\mathbb{C}[X,Y]$.
Proof.
Suppose $K=AB$ for two non-constant polynomials $A,B\in\mathbb{C}[X,Y]$. Given the degree of $K$, the polynomials $A$ and $B$ must be of degree $1$:
$A(x,y):=a_{1}x+a_{2}y+a_{3},\quad\quad B(x,y):=b_{1}x+b_{2}y+b_{3}$ (18)
where $a_{2}b_{2}=\sigma_{22}/2\neq 0$. Solving equation $K(x,y)=A(x,y)B(x,y)=0$ in $y$ yields
$\{Y^{+}(x),Y^{-}(x)\}=\{-(a_{1}x+a_{3})/a_{2},-(b_{1}x+b_{3})/b_{2}\}.$
This means that $Y^{+}$ and $Y^{-}$ are affine functions which implies that the polynomial under the square root of Equation (16) is the square of an affine function. This is equivalent to say that
$(\sigma_{12}x+\mu_{2})^{2}-\sigma_{22}(\sigma_{11}x^{2}+2\mu_{1}x)$
has a double root and then to the fact that the discriminant
$(\mu_{2}\sigma_{12}-\mu_{1}\sigma_{22})^{2}+\det(\Sigma)\mu_{2}^{2}$
(already computed in (17)) is equal to 00. This is not possible because $\det\Sigma>0$ since $\Sigma$ is a positive-definite covariance matrix. Hence there is no such decomposition of $K$, and $K$ is therefore irreducible.
We can sketch an alternative proof. According to the classification of conics (see [22], p.63), the nature of the conic $\{(x,y)\in\mathbb{R}^{2}:K(x,y)=0\}$ only depends on the signs of $\det(\Sigma)$ and the following block matrix determinant
[…]
The matrix $\Sigma$ is positive-definite, and so is its inverse $\Sigma^{-1}$, hence $\det(\Sigma)>0$ and $\Delta<0$, which corresponds to a non-degenerate ellipse. If $K=AB$ for two polynomials of degree $1$, it would imply that the ellipse would be equal to the union of two lines and therefore would be degenerate. We deduce again that $K$ is irreducible.
Note that the discriminant of the first proof is linked to the determinant of the second proof since $(\mu_{2}\sigma_{12}-\mu_{1}\sigma_{22})^{2}+\det(\Sigma)\mu_{2}^{2}=-\sigma_{22}\Delta$. ∎
It will be handy to work with the following rational uniformization of $\mathcal{S}$, first stated in [11, Proposition 5], defined by
@@
Figure 3: On the left, the points introduced in Lemmas 23 and 14 on the unit complex circle $\mathbb{U}$; on the right, their images by $(\bm{\mathrm{x}},\bm{\mathrm{y}})$ on the ellipse $\mathcal{S}\cap\mathbb{R}^{2}$. For the sake of readability, points on the ellipse are labelled with their preimages. Invariances in (22) are represented by the dotted arrows.
@@
    $G(y)=\frac{k_{1}(X^{+}(y),y)}{k_{2}(X^{+}(y),y)}\frac{k_{2}(X^{+}(y),\overline{y})}{k_{1}(X^{+}(y),\overline{y})}.$ (29)
@@ 3.1 Decoupling and Tutte’s invariant
Equation (32) justifies the terminology conformal gluing function for $w$, often encountered in the literature on Tutte’s invariant method.
Lemma 5.3 of [5] also provides, for some constant $\hat{C}$, the asymptotics
@@ 3.2 Explicit expression for the Laplace transforms
We now state our first main result when $\alpha\in\mathbb{N}$.
Theorem 10 (Laplace transforms, $\alpha\in\mathbb{N}$).
If $\alpha\in\mathbb{N}$ then the rational function defined by
$L(x,y):=\frac{k_{1}(x,y)P(x)+k_{2}(x,y)Q(y)}{K(x,y)}$ (42)
@@
Assuming that $\alpha=d-1\in\mathbb{N}$, the polynomials $P$ and $Q$ given in (34) form a decoupling pair by Proposition 8. The boundary value problem of Proposition 4 thus implies that $Q\varphi_{1}$ is an invariant, see (31). According to Lemma 7, we only need to prove that $Q\varphi_{1}$ doesn’t have any pole on $\overline{\mathcal{G}}_{\mathcal{H}}$, and has a finite limit as $y$ goes to +$\infty$+ ∞. By (27) and (34), and since $\alpha=d-1$ we have
@@
Comparing with Definition 30 we can see that the constant $\lambda$ given in (40) is equal to $-1$. The polynomial $k_{2}Q+k_{1}P$ vanishes on $\mathcal{S}$ the set of the zeros of $K$. By Hilbert’s Nullstellensatz [18],
@@
where $(K):=\{LK,L\in\mathbb{C}[X,Y]\}\text{ and }\sqrt{(K)}:=\{H\in\mathbb{C}[X,Y]:\exists m\in\mathbb{N}\text{ s.t. }H^{m}\in(K)\}$ are the ideal generated by $K$ and its radical. By Lemma 2, $K$ is irreducible which implies that $\sqrt{(K)}=(K)$. This shows that there exists $L\in\mathbb{C}[X,Y]$ such that $L(x,y)K(x,y)=k_{1}(x,y)P(x)+k_{2}(x,y)Q(y).$ We see that $L(x,y)\in\mathbb{R}$ for all $(x,y)\in\mathbb{R}^{2}$, so the coefficients of $L$ must be real. Substituting the values for $\varphi_{1}$ and $\varphi_{2}$ into the functional equation (13) yields $\varphi(x,y)=L(x,y)/(P(x)Q(y))$. ∎
@@
$S(z):=\displaystyle\prod_{q\in\mathrm{Z}}\frac{z-w(q)}{w(0)-w(q)},\text{ if }d\geqslant 2\quad\text{and}\quad S(z):=\displaystyle\prod_{p\in\mathrm{P}}\frac{w(0)-w(p)}{z-w(p)},\text{ if }d\leqslant 0.$ (47)
We deduce sufficient conditions for $\varphi_{1}$, $\varphi_{2}$ and $\varphi$ to belongs to the hierarchy (6). If $\alpha\in\mathbb{N}$ these Laplace transforms are rational, if $\pi/\beta\in\mathbb{Q}$ and $\alpha\in\mathbb{Z}+\frac{\pi}{\beta}\mathbb{Z}$ they are algebraic, if $\alpha\in\mathbb{N}+\frac{\pi}{\beta}\mathbb{Z}$ they are D-finite and if $\alpha\in\mathbb{Z}+\frac{\pi}{\beta}\mathbb{Z}$ they are D-algebraic.